Find the distance of the point (3,4,-4) from the line r(t) = (1 + 3t, -1 + 2t, 5 – 3t). Answer:
Answer #1

đxh = -3î +217 +11 fe .. magnitude taxi) = √3)²+(21)²+(192 - √9+4414 121 11@xB1 - 5571 (a)=137 +23 -3) lal = √9+4+9 = √22 lal